Suppose the price ratio of two goods is 3/4 and Shubman has a budget of $100. If his budget increases to $150 and the prices of the two goods stay the same, what is the new price ratio?


1
Expert's answer
2021-10-10T16:31:45-0400

Price ratio"=\frac{3}{4}"

Budget increases from $100 to $150.

"\implies" Budget increases by $50 which is half the original budget.

Thus, price ratio will also increase by half the original ratio:

"\frac{3}{4}\times\frac{1}{2}=\frac{3}{8}"

The new price ratio will be:

"\frac{3}{4}+\frac{3}{8}=\frac{9}{8}" .

New ratio is "\frac{9}{8}" .